OpenStudy (anonymous):

find the percent of increase. round to the nearest tenth, if necessary. from 45 m to 54 m a. 500% increase c. 20% increase b. 16.7% d. 83.3%

OpenStudy (akashdeepdeb):

Let the percent increase be = x% Thus 54 is 45 + x% of 45 [Because it is increasing by that percent] So\[\frac{x}{100} * 45 + 45 = 54\]

OpenStudy (akashdeepdeb):

Now can you try it? :)

OpenStudy (anonymous):

\[\frac{ 54-45 }{45 }*100=?\]

OpenStudy (anonymous):

-46?

OpenStudy (anonymous):

no

OpenStudy (akashdeepdeb):

That maybe the answer but does that seem logical to you? :D

OpenStudy (anonymous):

oops wait, would it be 20

OpenStudy (akashdeepdeb):

Can -46 be an INCREASE? :D

OpenStudy (akashdeepdeb):

Yes, correct. :D

OpenStudy (anonymous):

sorry, my calculator went wacko

OpenStudy (anonymous):

so the answer is 20? .-. :o

OpenStudy (akashdeepdeb):

Yes. Did you understand the concept though?

OpenStudy (anonymous):

yes, thank you so much :) it will definitely help me out with the other questions that are like this one now. I was having trouble with these types of questions so i posted an example. Thanks again! :D

OpenStudy (akashdeepdeb):

You're welcome. And also Welcome to OpenStudy. :)

OpenStudy (anonymous):

Thanks :D